Given consumption = 100 +0.75Yd

Tax = 50 + 0.5Y

Export = 200

Import = 50 + 0.25Y

Government spending = 150

Investment = 200

(a) Determine the value of the economy’s multiplier, which is applicable to government spending, and interpret it. 


1
Expert's answer
2021-05-13T17:55:55-0400

The disposable income Yd is;

Yd=Y-Taxes=Y-50-0.5Y

Yd=0.5Y-50

In equilibrium;Y=C+I+G+X-M

Y=100 +0.75(0.5Y-50)+I+G+X-50-0.25Y

Y=50+0.375Y-37.5+I+G+X-0.25Y

(1-0.375+0.25)Y=12.5+I+G+X

0.875Y=12.5+I+G+X

Y=1.14(125+I+G+X)

The government spending (G) multiplier in the economy is 1.14. It means that when G rises by $1, Y will rise by $1.14.